Let A = {2, 3, 6, 7} and B = {7, 3, -2, -8} FIND.
i. A∪B
ii. B∩A
iii. n(A) + n(B)
iv. n(A∪B)

Answer by Fombitz(32388)


i. A∪B={-8,-2,2,3,6,7}
ii. B∩A={3,7}
iii. n(A) + n(B)=4+4=8
iv. n(A∪B)=6